OpenStudy (anonymous):

a recipe for a soup calls for 2/4 cup of chopped onion and 1/5 cup of chopped celery .What is the total amount of celery and onion needed for the soup ?

OpenStudy (anonymous):

aren't you leaving some info out?

OpenStudy (anonymous):

or so you simply want to add 2/4 and 1/5?

OpenStudy (anonymous):

2/4 + 1/5 = 10/20 + 4/20 = 14/20 = 7/10